Three years ago Elaine was five times as old as Jerry. Today she is three times as old as Jerry. How old are Jerry and Elaine today?
Answer by stanbon(75887)

Three years ago Elaine was five times as old as Jerry. Today she is three times as old as Jerry. How old are Jerry and Elaine today?
----------
Equations:
e-3 = 5(j-3)
e = 3j
--------
Substitute for "e" and solve for "j":
3j-3 = 5j-15
2j = 12
Jerry = 6 yrs (Jerry[s age now)
-----------------------------------------
Solve for "e":
e = 3j = 18 yrs (Elaine's age now)
========================
Cheers,
Stan H.